# Prismatic-Bending Transformable (PBT) Joint for a Modular, Foldable Manipulator with Enhanced Reachability and Dexterity

### v1

@@ VI Conclusion and Future Work
In this paper, we introduced the Prismatic-Bending Transformable (PBT) Joint, a novel robotic joint designed to enhance adaptability, dexterity, and manipulability in modular robotic systems for dynamic and unstructured environments. The PBT Joint seamlessly integrates multi-directional bending, rotation, and elongation/contraction capabilities, supported by a direction-maintenance mechanism to ensure stability and proper configuration. We demonstrated two primary applications: a single-joint PBT wrist extension for existing manipulators, and a two-joint PBT modular manipulator capable of foldable and reconfigurable configurations. Experimental results validated the PBT Joint’s ability to improve workspace coverage, enhance dexterity in confined and clustered environments, and achieve enhanced obstacle avoidance.
Future work on the PBT joint and its modular manipulator includes several promising directions to advance their capabilities. Configurations with multiple interconnected modules could address more complex tasks, enhancing the versatility of modular manipulators. Miniaturizing the PBT joint to a centimeter scale or smaller would enable fine manipulations, such as finger-level tasks or micro-robotics. Integrating PBT-based systems into advanced platforms, including humanoid robots and autonomous manipulators, could significantly expand their operational scope. Additionally, improving the design to support higher bending payloads—currently limited compared to linear payloads—and enhancing energy efficiency would broaden their industrial applications, unlocking new potential.

### v2

@@ VI Conclusion and Future Work
In this paper, we introduced the Prismatic-Bending Transformable (PBT) Joint, a novel robotic joint that integrates multi-directional bending, rotation, and elongation/contraction within a compact, modular, and reconfigurable architecture. Inspired by scissor mechanisms and designed with directional maintenance consideration, the PBT Joint addresses limitations of conventional joint-link chain manipulators—particularly the lack of integrated dexterity and structural transformability. We demonstrated its versatility through two representative applications: a single-joint PBT wrist extension for existing robotic arms and a dual-joint foldable modular manipulator. Both configurations exhibited enhanced dexterity, expanded workspace coverage, and improved obstacle avoidance in confined and cluttered environments. Comprehensive validation through detailed design, mechanical optimization, kinematic and dynamic analysis, and experimental testing confirmed the PBT Joint’s effectiveness in enabling adaptable manipulation across a broad range of task scenarios.
Future work will focus on developing advanced control and motion planning strategies tailored to the PBT joint’s unique transformable capabilities, enabling precise, adaptive manipulation in cluttered and dynamic environments. Additional directions include scaling to multi-joint configurations, miniaturization for fine manipulation, and integration with humanoid and autonomous systems. Enhancing payload capacity and energy efficiency will further support deployment in demanding industrial applications.